Find the critical angle for an air–water boundary. (The index of refraction of water is 1.33.)


1
Expert's answer
2021-07-08T10:58:58-0400

From Snell's law we have,


a"\mu"w = "\frac{1}{Sin(ic)}"


Sin(ic) = "\frac{1}{aμw}"


= "\frac{1}{1.33}"


ic = 48o45'